OpenStudy (anonymous):

(3w-6/9) = (3) - (w-3/3)

OpenStudy (anonymous):

(3w - 6/9) = 3 - (w - 3/3) 3w - 2/3 = 3 - w + 1 3w - 2/3 = 4 - w 4w = 14/3 w = 14/2

OpenStudy (anonymous):